A circle is inscribed in a square with 10-inch sides. Find the area of the square not covered by the circle to the nearest tenth.
Answer by Edwin McCravy(20059)

A circle is inscribed in a square with 10-inch sides. Find the area of the square not covered by the circle to the nearest tenth.


Area of the square MINUS the area of the circle.

The side of the square is 10 inches, which is also the diameter of the circle
The radius of the circle is one-half the diameter, so r = 5 inches

Area of square = side² = 10² = 100 square inches
Are of circle = pr² = 3.14×5² = 3.1416×25 = 78.54

Area of the square MINUS the area of the circle = 100 - 78.54 = 21.46 sq. in.

To the nearest tenth, 21.5 square inches

Edwin